OpenStudy (anonymous):

I need an example of a function for which the domain is the set of all real numbers.

OpenStudy (anonymous):

f(x)=x

OpenStudy (anonymous):

you should tell whoever asked this question that the domain of function is supposed to be part of the definition, not something you find!

OpenStudy (anonymous):

just an example.

OpenStudy (anonymous):

but of course gianfranco is correct. you could say \[f(x)=x\] \[f(x)=5\] \[f(x)=x^2\]

OpenStudy (anonymous):

Thank you sat73 for you acknowledgenment, you are also correct in your responsew but he only asked for an example.

OpenStudy (anonymous):

just one of my silly pet peeves. "find the domain" is nonsense. your function should come equipped with the domain as part of its definition

OpenStudy (anonymous):

True. Enjoy my rarely given medal:)

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

Thisw was the way the instructor asked the question

OpenStudy (anonymous):

Agreed with satellite, it's nonsensical to ask for a domain independent of the function

OpenStudy (anonymous):

Business Manager, do not worry. You have all you need for the answer asked, here.

OpenStudy (anonymous):

tell you instructor that he/she doesn't understand the definition of a function. ok don't do that. don't even bring it up. the idea some people have is that if no domain is specified you are supposed to take it as the largest possible set of numbers for which the output would be a number. in practice this means you are not allowed to divide by zero or take an even root of a negative number. so for example the "domain" of \[f(x)=\sqrt{x}\] would be \[x\geq 0\] while the domain of \[f(x)=x\] would be all real numbers

OpenStudy (anonymous):

i will be quite now

OpenStudy (anonymous):

Alrighty then thanks for explaining